# Shannon Capacity of Channels with Markov Insertions, Deletions and Substitutions

### v1

@@ 5.2 Main Theorems
where $w={\widetilde{f}}_{n}(x)$ is the input vector $x$ without $l$-blocks and the last incomplete block, i.e. ${\widetilde{f}}_{n}(x)=(x^{(1)},x^{(2)},\dots,x^{(q)})$ , $x^{(i)}=x_{i(m+l)-m+1}^{i(m+l)}$. Denote by ${\overline{p}}_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n^{\prime}}}$, $n^{\prime}=qm$ the marginalized distribution of ${\widetilde{f}}_{n}(x)$, given distribution $p_{n}$ over $x$:
@@
The fact that channel probabilities of ${\widetilde{V}}_{n}[\rho]$ and $V^{q}_{m}[\pi]$ are close implies the fact that mutual information density of ${\widetilde{V}}_{n}[\rho]$ and $V^{q}_{m}[\pi]$ are close too, since
@@
$\lim_{n\to\infty}{\mathbf{E}}{\left[{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)-\mathbf{i}(V^{q}_{m}[\pi],{\overline{p}}_{n},x,y_{\text{even}})}\right]}$
$=\lim_{n\to\infty}{\mathbf{E}}{\left[{\left|{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)-\mathbf{i}(V^{q}_{m}[\pi],{\overline{p}}_{n},x,y_{\text{even}})}\right|}\right]}=0.$
@@
$=\lim_{n\to\infty}{\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)}{n}-\frac{q}{n}\cdot{\mathbf{I}}(V_{m}[\pi])}\right|}\right]}=\lim_{n\to\infty}{\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|}\right]}=0$ (47)
By (46), $\lim_{n\to\infty}{\mathbf{I}}({\widetilde{V}}_{n}[\rho],p_{n})/n={\mathbf{I}}(V)$, so $p_{n}$ allows channel sequence ${\widetilde{V}}_{n}[\rho]$ to have i-capacity equal to ${\mathbf{I}}(V)$ which means ${\mathbf{I}}({\widetilde{V}}_{n}[\rho])\geq{\mathbf{I}}(V)$. It can be shown [^5] Note that ${\widetilde{V}}_{n}[\rho]$ is obtained from $V^{q}_{m}[\pi]$ by deleting $o(n)$ component channel outputs, each of which has entropy upper-bounded by ${\overline{H}}$. Furthermore, inputs of $V_{m}^{q}[\pi]$ independently influence the output, so no distribution $p$ for ${\widetilde{V}}_{n}[\rho]$ can lead to ${\widetilde{p}}$ which has strictly larger m.i. capacity ${\mathbf{I}}(V^{q}_{m}[\pi],{\widetilde{p}})$, than independent maximization over distributions of each $x^{(i)}$. That said, we cannot lose more than $o(n)\cdot{\overline{H}}$ bits of information by transition from ${\widetilde{V}}_{n}[\rho]$ to $V^{q}_{m}[\pi]$. that ${\mathbf{I}}({\widetilde{V}}[\rho])={\mathbf{I}}(V)$. Thus, (47) implies
$\lim_{n\to\infty}{\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)}{n}-{\mathbf{I}}({\widetilde{V}}_{n}[\rho])}\right|}\right]}=0,$
which is exactly the information stability condition for ${\widetilde{V}}[\rho]$, which by Proposition 1 implies that ${\mathbf{I}}({\widetilde{V}}[\rho])={\mathbf{C}}({\widetilde{V}}[\rho])$, and also, since ${\mathbf{I}}({\widetilde{V}}[\rho])={\mathbf{I}}(V)$, the m.i. capacity does not depend on $\rho$, so we can underline it by writing
${\mathbf{I}}({\widetilde{V}}[\rho])={\mathbf{C}}({\widetilde{V}}[\rho])={\mathbf{I}}({\widetilde{V}})={\mathbf{C}}({\widetilde{V}}).$ (48)
Recalling (45), one obtains ${\mathbf{C}}({\overline{V}})={\mathbf{I}}({\overline{V}})$, and by (44), ${\mathbf{C}}(V)={\mathbf{I}}(V)$.
6 Conclusion
In this paper, we have focused on channels with synchronization errors, for which the insertions/deletions are governed by a Markov chain. Specifically, we have shown that if the underlying Markov chain is stationary and ergodic, then the information capacity of the corresponding channel sequence exists and it is equal to the coding capacity. This result generalizes the classical result of Dobrushin on the existence of the Shannon capacity for channels with memoryless synchronization errors. To accomplish this goal, we generalized methods used in [2, 3] as separate independent propositions. These propositions state that one can apply a function sequence to a channel sequence, changing neither the coding capacity nor the mutual information capacity of the latter, if the function sequence has asymptotically small preimage or asymptotically small probability of changing the output. By applying such function sequences, one can bring the original channel sequence to another channel sequence, for which the capacity theorem or capacity bounds can be more easily derived, and one can be sure that the coding capacity / mutual information capacity is not changed under such transformation. The methodology may be useful for deriving capacity theorems and capacity bounds for other non-trivial channel sequences.

### v3

@@ V-B Main Theorems
where $w={\widetilde{f}}_{n}(x)$ is the input vector $x$ without the $l$-blocks and the last incomplete block, i.e. ${\widetilde{f}}_{n}(x)=(x^{(1)},x^{(2)},\dots,x^{(q)})$, $x^{(i)}=x_{i(m+l)-m+1}^{i(m+l)}$. Denote by ${\overline{p}}_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n^{\prime}}}$, $n^{\prime}=qm$ the marginalized distribution of ${\widetilde{f}}_{n}(x)$, given distribution $p_{n}$ over $x$:
@@
The fact that channel probabilities of ${\widetilde{V}}_{n}[\rho]$ and $V^{q}_{m}[\pi]$ are close implies that mutual information densities of ${\widetilde{V}}_{n}[\rho]$ and $V^{q}_{m}[\pi]$ are close too, since
@@
$\lim_{n\to\infty}{\mathbf{E}}{\left[{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)-\mathbf{i}(V^{q}_{m}[\pi],{\overline{p}}_{n},x,y_{\text{even}})}\right]}=\lim_{n\to\infty}{\mathbf{E}}{\left[{\left|{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)-\mathbf{i}(V^{q}_{m}[\pi],{\overline{p}}_{n},x,y_{\text{even}})}\right|}\right]}=0.$
@@
$=\lim_{n\to\infty}{\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)}{n}-\frac{q}{n}\cdot{\mathbf{I}}(V_{m}[\pi])}\right|}\right]}=\lim_{n\to\infty}{\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|}\right]}=0.$ (49)
By (46) and (47), ${\mathbf{I}}({\widetilde{V}}[\rho])={\mathbf{I}}(V)$. By (49), the channel sequence ${\widetilde{V}}[\rho]$ satisfies Proposition 2, which implies ${\mathbf{I}}{\mathbf{S}}({\widetilde{V}}[\rho])$, which by (46) and (47) is equivalent to ${\mathbf{I}}{\mathbf{S}}(V)$. ∎
VI Conclusion
In this paper, we have focused on channels with synchronization errors, for which the insertions/deletions are governed by a Markov chain. Specifically, we have shown that if the underlying Markov chain is stationary and ergodic, then the information capacity of the corresponding channel sequence exists and it is equal to the coding capacity. This result generalizes the classical result of Dobrushin on the existence of the Shannon capacity for channels with memoryless synchronization errors. To accomplish this goal, we generalized methods used in [4, 5] as separate independent propositions. These propositions state that one can apply a function sequence to a channel sequence, changing neither the coding capacity nor the mutual information capacity of the latter, if the function sequence has asymptotically small preimage or asymptotically small probability of changing the output. By applying such function sequences, one can bring the original channel sequence to another channel sequence, for which the capacity theorem or capacity bounds can be more easily derived, and one can be sure that the coding capacity / mutual information capacity is not changed under such transformation. The methodology may be useful for deriving capacity theorems and capacity bounds for other non-trivial channel sequences. The future work can be done towards generalization of this paper and [10] in order to include simultaneously the state’s randomness and its dependence on input symbols.